Check whether the following sequences are convergent:

1){4+(-1)^n}

2){(4n+n^2)/(2n^2+3n)}
1
Expert's answer
2020-03-09T13:38:54-0400

1) "x_n=4+(-1)^n"

"\lim\limits_{n\to\infty}x_n=\left
\{\begin{matrix}
 4+(-1)^{n}=5, n=2k \\
 4+(-1)^{n}=3, n=2k-1 
\end{matrix}\right."

because limit  takes two different values ​​then

the sequence is divergent

2) "x_n=\frac{4n+n^2}{2n^2+3n}"

consider the limit

"\lim\limits_{n\to\infty}x_n=\lim\limits_{n\to\infty}\frac{4n+n^2}{2n^2+3n}=\\
=\lim\limits_{n\to\infty}\frac{n^2(\frac{4}{n}+1)}{n^2(2+\frac{3}{n})}=\\
=\lim\limits_{n\to\infty}\frac{\frac{4}{n}+1}{2+\frac{3}{n}}=\frac{1}{2}"

the sequence is convergent